im have trouble with this ? in the back of my grade 10 math book the answer is 1/6 here is the question
-3/y +6(4)=6
-3/y +24= 6 so i bring the 24 over it become a - now
= 6-24 =18
= -18
for the -3/y do i divide that into the -18, i thought when u bring it over to the other side i have to times it -3/y * -18


Answer by pwac(253)

-3/y+24=6
-3/y=6-24
-3/y=-18
-3=-18y
-3/-18=y(-ve dividing -ve = +ve)
y=1/6
Pete